# On Selection of Cross-Section Averages in Non-stationary Environments

### v2

@@ 6.5 Proof Changes for IC of Margaritella and Westerlund, (2023)
$\widehat{\boldsymbol{\nu}}_{i}=\mathbf{y}_{i}-\mathbf{X}_{i}\widehat{\boldsymbol{\beta}}=-\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})+\mathbf{X}_{i}\boldsymbol{\upsilon}_{i}+\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i},$ (A.84)
@@
where now $\left\|\sqrt{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|=O_{p}(1)$ under general factors (see e.g. Stauskas,, 2023). This means that now, for example,
@@
$\left\|\frac{1}{NT}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|$ $\leq N^{-1}\underbrace{\left\|\sqrt{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|^{2}}_{O_{p}(1)}\underbrace{\left\|\frac{1}{NT}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}\right\|}_{O_{p}(T^{\tau})}$
=$=\begin{cases}O_{p}(N^{-1}T^{\tau}),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ o_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M,\end{cases}$ (A.85)
and other terms will behave similarly, while
@@
$\left\|\frac{1}{NT}\sum_{i=1}^{N}\boldsymbol{\upsilon}_{i}^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}\boldsymbol{\upsilon}_{i}\right\|=\begin{cases}O_{p}(T^{\tau}),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ o_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M,\end{cases}$ (A.86)
[…]
as before, whereas the cross-terms are of lower order that could be made negligible by imposing restrictions on $N$ and $T$ divergence rate.
@@ 6.6.1 Independent Idiosyncratics
Table 3: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$. DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.2) and (2.1). For $IC^{DVS}_{1}$ and $IC^{MW}_{1}$, $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $IC^{DVS}_{2}$ and $IC^{MW}_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$. Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$, and $\varepsilon_{i,t}$ are uncorrelated over time, but weakly correlated across units, see (3.2) in the main text.
@@
Table 4: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$ with generated factors, $\overline{\mathbf{Z}}_{M}=\mathbf{F}\mathbf{q}_{M}$. DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.1) and (2.2). For $DVS_{1}$ and $MW_{1}$, $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $DVS_{2}$ and $MW_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$. Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$, and $\varepsilon_{i,t}$ are uncorrelated over time, but weakly correlated across units, see (3.2) in the main text.

### v3

@@ 5.3.3.2  Heterogeneous $\boldsymbol{\beta}$
$\widehat{\boldsymbol{\nu}}_{i}=\mathbf{y}_{i}-\mathbf{X}_{i}\widehat{\boldsymbol{\beta}}$ $=-\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})+\mathbf{X}_{i}\boldsymbol{\upsilon}_{i}+\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}$
$=-\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})+\mathbf{F}\boldsymbol{\Gamma}_{i}\boldsymbol{\upsilon}_{i}+\mathbf{V}_{i}\boldsymbol{\upsilon}_{i}+\mathbf{F}\boldsymbol{\gamma}_{i}+\boldsymbol{\varepsilon}_{i}$
$=-\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})+\mathbf{F}(\boldsymbol{\Gamma}_{i}\boldsymbol{\upsilon}_{i}+\boldsymbol{\gamma}_{i})+\mathbf{V}_{i}\boldsymbol{\upsilon}_{i}+\boldsymbol{\varepsilon}_{i}$
$=-\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})+\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*},$ (A.95)
@@
with the obvious definitions of $\boldsymbol{\gamma}_{i}^{*}$ and $\boldsymbol{\varepsilon}_{i}^{*}$. Note that now $\left\|\sqrt{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|=O_{p}(1)$ under general factors (see e.g. Stauskas,, 2023), which is a considerably lower rate. This means that now, for example,
@@
$\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|$ $\leq N^{-1}\underbrace{\left\|\sqrt{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|^{2}}_{O_{p}(1)}\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\mathbf{X}_{i}\right\|$
$=\begin{cases}O_{p}(N^{-1}),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ o_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M,\end{cases}$ (A.96)
and other terms will behave similarly:
@@
$\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})^{\prime}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})\right\|$
$\leq N^{-1/2}\left\|\sqrt{N}(\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta})\right\|\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\mathbf{X}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})\right\|$
$=\begin{cases}O_{p}(N^{-1/2}),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ o_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M.\end{cases}$ (A.97)
Finally,
$\left\|\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})\right\|=\begin{cases}O_{p}(T^{\tau}),\hskip 2.84526pt\text{under}\hskip 2.84526ptM\subset M_{0}\\ O_{p}(C_{N,T}^{-2})=o_{p}(1),\hskip 2.84526pt\text{under}\hskip 2.84526ptM_{0}\subset M,\end{cases}$ (A.98)
and, in particular, under $M\subset M_{0}$:
$\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})$
$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})^{\prime}(\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}-\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M_{0}}})(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})+o_{p}(T^{-\tau})$
$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}(\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+\boldsymbol{\varepsilon}_{i}^{*})+o_{p}(T^{-\tau})$
$=\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\boldsymbol{\gamma}_{i}^{*\prime}\mathbf{F}^{\prime}\mathbf{P}_{\mathbf{M}_{\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M}}\mathbf{F}\overline{\boldsymbol{\Gamma}}\mathbf{q}_{M^{C}}}\mathbf{F}\boldsymbol{\gamma}_{i}^{*}+o_{p}(T^{-\tau})$
$=d_{NT}+o_{p}(T^{-\tau})$ (A.99)
with $d_{NT}\to_{p}d>0$ in analogy to the analysis in (5.3.3.1). Now, we need to assume $T^{\tau}N^{-1/2}\to 0$ in addition to $\tau\in(0,0.5)$ to isolate the $o_{p}(T^{-\tau})$ terms. Then in the under-specification case, we have
$\frac{1}{NT^{1+\tau}}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}(\mathbf{M}_{\widehat{\mathbf{F}}_{M}}-\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}})\widehat{\boldsymbol{\nu}}_{i}=d_{NT}+o_{p}(T^{-\tau}),$ (A.100)
and, since again we can show that $e_{NT}:=\frac{1}{NT}\sum_{i=1}^{N}\widehat{\boldsymbol{\nu}}_{i}^{\prime}\mathbf{M}_{\widehat{\mathbf{F}}_{M_{0}}}\widehat{\boldsymbol{\nu}}_{i}$ converges to a positive scalar quantity, we can demonstrate that
$(\mathrm{IC}^{MW}(M)$ $-\mathrm{IC}^{MW}(M_{0}))/\ln(T)=\ln(1+T^{\tau}(d_{NT}\times e_{NT}^{-1}+o_{p}(T^{-\tau}))/\ln(T)+o(1)$
$=\ln(1+T^{\tau}d_{NT}\times e_{NT}^{-1})/\ln(T)+o_{p}(1)$
$\to_{p}\tau>0$ (A.101)
as $(N,T)\to\infty$. Therefore,
$\mathbb{P}(\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0})<0)=\mathbb{P}\left(\frac{\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0})}{\ln(T)}<0\right)\to 0$ (A.102)
as $(N,T)\to\infty$. Under $M_{0}\subset M$, the proof once again the same as in Margaritella and Westerlund, (2023),and so given the rates declared above and the fact that $p_{N,T}^{-1}O_{p}(C_{N,T}^{-2})=o(1)$, we get
$p_{N,T}^{-1}(\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0}))=(g-m)+o_{p}(1)$ (A.103)
and so
$\mathbb{P}\left(\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0})<0\right)=\mathbb{P}\left(p_{N,T}^{-1}(\mathrm{IC}^{MW}(M)-\mathrm{IC}^{MW}(M_{0}))<0)\right)\to 0,$ (A.104)
as expected.
@@ 5.4.1 Independent Idiosyncratics
Stationary $\mathbf{F}$ Non-Stationary $\mathbf{F}$
$\tau=0$ $\tau=0.4$ $\tau=0.9$
NN $T$ $IC^{MW}_{1}$ $IC^{MW}_{2}$ $IC^{DVS}_{1}$ $IC^{DVS}_{2}$ $IC^{MW}_{1}$ $IC^{MW}_{2}$ $IC^{DVS}_{1}$ $IC^{DVS}_{2}$ $IC^{MW}_{1}$ $IC^{MW}_{2}$ $IC^{DVS}_{1}$ $IC^{DVS}_{2}$
Correct Selection Frequency for $g$
50 50 31.00 15.90 36.40 16.10 7.30 3.60 3.50 0.30 0.10 0.00 0.00 0.00
100 50 97.30 95.00 99.20 96.40 91.50 82.80 95.10 85.90 11.50 8.60 13.40 7.50
200 50 95.50 94.30 99.70 98.80 48.60 45.50 77.70 69.20 0.20 0.20 5.50 4.10
300 50 100.00 100.00 99.60 99.50 81.30 78.20 69.90 67.70 4.30 3.70 1.80 1.60
500 50 100.00 100.00 100.00 100.00 99.60 99.40 96.30 95.10 21.20 20.10 11.50 10.10
50 100 97.10 89.30 100.00 100.00 68.60 55.30 99.30 94.50 7.00 5.90 18.20 10.90
100 100 100.00 99.90 99.90 99.00 62.50 49.40 56.30 39.70 2.70 1.30 1.70 0.60
200 100 100.00 100.00 100.00 100.00 99.90 99.60 99.90 99.70 24.50 19.70 26.30 18.80
300 100 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 39.10 34.30 46.60 39.30
500 100 100.00 100.00 100.00 100.00 99.70 99.70 95.90 94.80 16.30 15.30 8.50 7.60
50 200 99.70 99.90 99.90 99.80 97.70 96.70 89.00 79.90 9.80 8.70 5.10 3.50
100 200 100.00 100.00 100.00 100.00 93.00 88.80 99.50 98.60 5.30 4.70 12.80 7.70
200 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 40.80 32.70 60.20 44.50
300 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 22.40 18.90 81.40 73.20
500 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 54.10 49.30 80.80 75.50
50 300 100.00 100.00 100.00 100.00 88.60 86.40 99.90 99.40 7.40 6.20 15.20 11.00
100 300 98.90 99.80 100.00 100.00 98.70 99.40 100.00 100.00 21.30 18.60 27.20 22.40
200 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 56.00 49.10 52.60 45.60
300 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 89.30 82.90 94.40 89.10
500 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 80.10 75.70 93.70 91.90
50 500 100.00 100.00 100.00 100.00 90.10 89.00 100.00 100.00 3.40 3.10 11.20 10.20
100 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 38.50 37.10 53.70 50.80
200 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 84.10 81.30 97.30 95.80
300 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 88.30 85.90 97.00 95.60
500 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 92.70 88.20 99.50 98.70
Average of $g$
50 50 3.16 2.48 3.14 1.96 2.41 2.02 2.27 1.68 1.51 1.38 1.40 1.23
100 50 3.97 3.94 3.99 3.96 3.88 3.71 3.95 3.82 2.11 1.96 2.23 1.96
200 50 3.96 3.94 3.99 3.98 3.43 3.38 3.56 3.38 1.86 1.82 1.61 1.54
300 50 4.00 4.00 4.00 4.00 3.80 3.77 3.70 3.67 2.21 2.17 2.14 2.07
500 50 4.00 4.00 4.00 4.00 4.00 3.99 3.96 3.95 2.65 2.61 2.37 2.32
50 100 3.96 3.86 4.00 4.00 3.65 3.34 3.99 3.92 2.08 1.95 2.31 2.03
100 100 4.00 4.00 4.00 3.99 3.45 3.19 3.51 3.21 1.45 1.33 1.73 1.52
200 100 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 2.69 2.54 2.62 2.41
300 100 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.14 3.05 3.21 3.06
500 100 4.00 4.00 4.00 4.00 4.00 4.00 3.96 3.95 2.63 2.58 2.76 2.70
50 200 4.00 4.00 4.00 4.00 3.98 3.96 3.84 3.70 1.93 1.86 1.58 1.50
100 200 4.00 4.00 4.00 4.00 3.92 3.87 3.99 3.98 2.09 2.02 1.95 1.77
200 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.16 2.98 3.36 3.04
300 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.01 2.92 3.61 3.44
500 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.21 3.11 3.69 3.56
50 300 4.00 4.00 4.00 4.00 3.89 3.86 4.00 3.99 2.14 2.09 2.24 2.11
100 300 4.01 4.00 4.00 4.00 4.01 4.00 4.00 4.00 2.83 2.75 2.69 2.57
200 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.50 3.38 3.48 3.38
300 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.85 3.74 3.94 3.87
500 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.80 3.75 3.93 3.91
50 500 4.00 4.00 4.00 4.00 3.90 3.89 4.00 4.00 1.92 1.89 1.98 1.94
100 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.08 3.04 3.29 3.22
200 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.81 3.77 3.97 3.95
300 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.87 3.83 3.97 3.96
500 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.90 3.83 3.99 3.98
@@
Table 2: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$. DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.1) and (2.2) in the main text. For $DVS_{1}$ and $MW_{1}$, $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $DVS_{2}$ and $MW_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$. Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$, and $\varepsilon_{i,t}$ are uncorrelated over time, but weakly correlated across units, see (3.2) in the main text.
Stationary $\mathbf{F}$ Non-Stationary $\mathbf{F}$
$\tau=0$ $\tau=0.4$ $\tau=0.9$
NN $T$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$
Correct Selection Frequency for $g$
50 50 82.60 74.70 37.60 16.30 26.30 17.90 2.60 0.40 0.10 0.00 0.00 0.00
100 50 100.00 100.00 99.20 95.20 99.90 99.70 94.10 83.60 42.00 33.80 7.20 4.30
200 50 100.00 100.00 99.60 98.70 99.90 99.90 75.70 67.40 28.20 24.80 2.30 1.40
300 50 100.00 100.00 99.60 99.50 93.90 92.80 69.70 67.00 8.80 7.40 1.50 1.10
500 50 100.00 100.00 100.00 100.00 100.00 100.00 96.20 94.90 49.80 47.70 9.70 8.50
50 100 100.00 100.00 100.00 99.40 100.00 100.00 97.20 89.80 42.40 32.10 7.50 3.90
100 100 100.00 100.00 99.90 98.40 90.50 79.90 47.90 31.80 2.80 1.70 0.40 0.20
200 100 100.00 100.00 100.00 100.00 100.00 100.00 99.90 99.70 78.90 72.30 19.30 12.90
300 100 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 85.50 83.20 41.80 36.10
500 100 100.00 100.00 100.00 100.00 100.00 99.80 96.00 94.70 27.10 25.20 7.40 6.80
50 200 100.00 100.00 100.00 99.60 100.00 100.00 85.50 73.00 59.30 53.30 0.60 0.50
100 200 100.00 100.00 100.00 100.00 100.00 100.00 99.50 98.40 49.80 42.90 5.50 3.80
200 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 95.20 92.20 49.70 36.30
300 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 98.80 96.90 77.10 67.70
500 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.80 99.70 76.40 69.70
50 300 100.00 100.00 100.00 100.00 100.00 100.00 99.50 98.90 58.50 55.40 4.80 4.00
100 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.90 85.50 82.80 16.70 12.50
200 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 91.50 88.50 48.10 40.50
300 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.90 99.70 91.40 85.60
500 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.90 99.70 93.10 90.90
50 500 100.00 100.00 100.00 100.00 100.00 100.00 99.80 99.80 17.40 16.00 3.00 2.30
100 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 95.70 95.00 40.70 37.10
200 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.90 99.90 95.70 94.20
300 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 95.80 94.20
500 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.30 97.80
Average of $g$
50 50 3.83 3.74 3.07 1.87 3.15 2.97 2.19 1.63 1.85 1.71 1.32 1.20
100 50 4.00 4.00 3.99 3.94 4.00 4.00 3.93 3.78 3.11 2.90 2.02 1.80
200 50 4.00 4.00 3.99 3.97 4.00 4.00 3.52 3.35 2.48 2.38 1.58 1.52
300 50 4.00 4.00 4.00 4.00 3.94 3.93 3.70 3.67 2.82 2.77 2.09 2.03
500 50 4.00 4.00 4.00 4.00 4.00 4.00 3.96 3.95 3.38 3.35 2.33 2.27
50 100 4.00 4.00 4.00 3.99 4.00 4.00 3.97 3.85 3.02 2.77 2.07 1.86
100 100 4.00 4.00 4.00 3.98 3.90 3.79 3.40 3.07 1.85 1.70 1.45 1.33
200 100 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.69 3.58 2.48 2.27
300 100 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.84 3.81 3.15 3.01
500 100 4.00 4.00 4.00 4.00 4.00 4.00 3.96 3.95 3.20 3.17 2.69 2.63
50 200 4.00 4.00 4.00 4.00 4.00 4.00 3.77 3.55 3.20 3.08 1.45 1.41
100 200 4.00 4.00 4.00 4.00 4.00 4.00 3.99 3.98 3.10 2.93 1.82 1.71
200 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.94 3.89 3.19 2.90
300 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.98 3.94 3.54 3.35
500 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.62 3.48
50 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.99 3.40 3.33 1.96 1.90
100 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.78 3.73 2.53 2.42
200 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.92 3.88 3.43 3.32
300 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.91 3.83
500 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.92 3.89
50 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 2.42 2.38 1.75 1.71
100 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.95 3.94 3.04 2.95
200 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.95 3.93
300 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.96 3.94
500 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.99 3.97
@@
Table 3: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$ with generated factors, $\overline{\mathbf{Z}}_{M}=\mathbf{F}\mathbf{q}_{M}$. DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.1) and (2.2) in the main text. For $DVS_{1}$ and $MW_{1}$, $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $DVS_{2}$ and $MW_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$. Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$, and $\varepsilon_{i,t}$ are uncorrelated over time, but weakly correlated across units, see (3.2) in the main text.